Determine the acceleration of the 10kg mass with all of the following forces acting on it, and express it in polar coordinates.
Force A: Fappx=204.79N Fappy= 143.39N
Force B: Fappx: -51.30N Fappy=140.95N
Force C: Fappx=117.46N Fappy=-42.75N
1
Expert's answer
2018-11-12T15:56:10-0500

a_x=(204.79-51.30+117.46 )/10=27.095 m/s^2 .

a_y=(143.39+140.95-42.75)/10=24.159 m/s^2 .

In polar coordinates:

a=√(〖24.159〗^2+(27.095)^2 )=36.302 m/s^2 .

θ=tan^(-1)⁡〖24.159/27.095〗=41.72°